# The Dynamic, the Static, and the Weak factor models and the analysis of high-dimensional time series

### v1

@@ 3.2 The General Dynamic Factor Model
It is important to stress that neither Forni and Lippi (2001) nor Hallin and Lippi (2013) impose any rate on the divergence, as $n\to\infty$, of the eigenvalues of the spectral density matrix ${\bm{\Sigma}}^{(n)}_{\bf X}$. Indeed, no divergence rate assumption is needed for identifying the model as long as a diverging gap exists between the exploding and bounded eigenvalues. When turning to estimation, Forni et al. (2000) still do not impose any rates and, because of this, cannot establish any consistency rates—on this latter issue, see Forni et al. (2004). Hallin and Liška (2007), Forni et al. (2017), and, in a more general context nesting the GDFM, Barigozzi et al. (2023), require the additional assumption of linear divergence rates, thus deriving also consistency rates for the estimators defined therein. This issue is discussed in Section 5.
4 The Approximate Static Factor Model revisited
@@
Paralleling the characterization (3.3) of the GDFM in the previous section, a representation-based characterization of the approximate static factor decomposition of Chamberlain and Rothschild also can be obtained, which extends (2.1).
@@
Denoting by ${\mathcal{H}}^{{\bf X}_{t}}$ the Hilbert space spanned by ${\bf X}_{t}$, consider the following decomposition of ${\mathcal{H}}^{{\bf X}_{t}}$ into a statically (at time $t$) common subspace ${\mathcal{H}}^{{\bf X}_{t}}_{\scriptscriptstyle\text{\rm stat com}}$ and its orthogonal complement (at time $t$ and within ${\mathcal{H}}^{{\bf X}_{t}}$) ${\mathcal{H}}^{{\bf X}_{t}}_{\scriptscriptstyle\text{\rm stat idio}}\coloneqq({\mathcal{H}}^{{\bf X}_{t}}_{\scriptscriptstyle\text{\rm stat com}})^{\perp}$.
@@
A random variable $\zeta$ with values in ${\mathcal{H}}^{{\bf X}_{t}}$ and variance $\sigma^{2}_{\zeta}$ is called statically common at time $t$ if either $\sigma^{2}_{\zeta}=0$ (hence $\zeta=0$ a.s.) or $\sigma^{2}_{\zeta}>0$ and $\zeta/\sigma_{\zeta}$ is the limit in quadratic mean, as $n\to\infty$, of a sequence of standardized elements of ${\mathcal{H}}^{{\bf X}_{t}}$ of the form ${w_{\bf X}^{(n)}}/\sigma^{(n)}_{w}$, where
@@
The spaces $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ and $\mathcal{H}_{\scriptscriptstyle\text{\rm stat idio}}^{{\bf X}_{t}}$ always exist and are unique for all $t$. Intuitively, $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ is the Hilbert space generated by the exploding (as $n\to\infty$) normed linear combinations of the $X_{it}$’s, $i=1,\ldots,n$, for given $t$. Contrary to its dynamic counterpart, it does depend on $t$. It readily follows from the definitions that, for all $t\in\mathbb{Z}$, $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}\subseteq\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{\bf X}$.
@@
where $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$ are the projections of $X_{it}$ on $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ and $\mathcal{H}_{\scriptscriptstyle\text{\rm stat idio}}^{{\bf X}_{t}}$, respectively. Orthogonality between the common and the idiosyncratic components $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$, however, only holds at time $t$—more precisely, $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{i^{\prime}t}^{\scriptscriptstyle\text{\rm stat}}$ are mutually orthogonal for all $i,\,i^{\prime}$, and $t$—while for $\chi_{it}^{\scriptscriptstyle\text{\rm dyn}}$ and $\xi_{it}^{\scriptscriptstyle\text{\rm dyn}}$ in (3.5) orthogonality of $\chi_{it}^{\scriptscriptstyle\text{\rm dyn}}$ and $\xi_{i^{\prime}t^{\prime}}^{\scriptscriptstyle\text{\rm dyn}}$ holds for all $i,\,i^{\prime}$, $t$, and $t^{\prime}$.
So far, (4.2) still is more general than (2.1). Adding the requirement that the eigenvalues
@@
of the $n\times n$ covariance matrices ${\bm{\Gamma}}^{(n)}_{{X}}$ of ${\bf X}^{(n)}_{t}\coloneqq(X_{1t},\ldots,X_{nt})^{\prime}$ (which, in view of second-order stationarity, does not depend on $t$) are such that
$\lim_{n\to\infty}\lambda^{(n)}_{X;r}=\infty\quad\text{ and }\quad\lim_{n\to\infty}\lambda^{(n)}_{X;r+1}<\infty,$ (4.3)
@@
for some finite $r\in{\mathbb{N}}$ independent of $n$, it can be shown (see the Appendix) that $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ admits $r$-dimensional orthonormal bases ${\bf f}_{t}=(f_{1t},\ldots,f_{rt})^{\prime}$, any of which can be interpreted as an $r$-tuple of factors, yielding a Chamberlain and Rothschild approximate static factor decomposition (2.1) satisfying the assumptions (a)–(g) of Section 2.1, with statically common component $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ of the form ${\bf B}_{i}{\bf f}_{t}$. This result is new to the literature and makes the static approximate factor model decomposition also a representation result rather than a statistical model.
@@
Note that, due to second-order stationarity, and despite the fact that $\mathcal{H}^{{\bf X}_{t}}$, $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$, and $\mathcal{H}_{\scriptscriptstyle\text{\rm stat idio}}^{{\bf X}_{t}}$ depend on $t$, ${\bm{\Gamma}}^{(n)}_{X}$, its eigenvalues and eigenvectors do not; accordingly, the number $r$ of diverging eigenvalues in (4.3) (which, under (4.3), is the number of factors and the dimension of $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$), their values and the corresponding eigenvectors, and the projection operators $\text{\rm Proj}_{\scriptscriptstyle\text{\rm com}}$ and $\text{\rm Proj}_{\scriptscriptstyle\text{\rm idio}}$ from $\mathcal{H}^{{\bf X}_{t}}$ to $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ and $\mathcal{H}_{\scriptscriptstyle\text{\rm stat idio}}^{{\bf X}_{t}}$ (which only depend on ${\bm{\Gamma}}^{(n)}_{X}$), are the same for all $t$.
@@
Finally, it is important to notice that Chamberlain and Rothschild (1983) do not impose any rate on the divergence, as $n\to\infty$, of the eigenvalues of the covariance matrix ${\bm{\Gamma}}^{(n)}_{\chi}$, which in principle can be sublinear, linear, or superlinear. Indeed, no divergence rate assumption is needed for identifying the model as long as an eigen-gap widening as $n\to\infty$, exists. When turning to estimation all works mentioned above make the additional assumption of a linear rate of divergence of the eigenvalues of the covariance matrix ${\bm{\Gamma}}^{(n)}_{\chi}$ (Stock and Watson, 2002a, Bai, 2003, Forni et al., 2009, etc.). The same linear divergence assumption is made by Bai and Ng (2002) when proposing an information criterion to determine $r$ based on the behavior of sample eigenvalues.

### v3

@@ 3.2 The General Dynamic Factor Model
It is important to stress that neither Forni and Lippi (2001) nor Hallin and Lippi (2013) impose any rate on the divergence, as $n\to\infty$, of the eigenvalues of the spectral density matrix ${\bm{\Sigma}}^{(n)}_{X}$. Indeed, no divergence rate assumption is needed for identifying the model as long as a diverging gap exists between the exploding and bounded eigenvalues. When turning to estimation, Forni et al. (2000) still do not impose any rates and, because of this, cannot establish any consistency rates—on this latter issue, see Forni et al. (2004). Hallin and Liška (2007), Forni et al. (2017), and, in a more general context nesting the GDFM, Barigozzi et al. (2024), require the additional assumption of linear divergence rates, thus deriving also consistency rates for the estimators defined therein. This issue is discussed in Section 6.
4 Orthogonality
A central issue in the theory and practice of factor models is orthogonality between the common and idiosyncratic components. The GDFM literature always naturally assumes that orthogonality between dynamically common and idiosyncratic components holds at all leads and lags, i.e., ${\rm E}[\chi_{it}^{\scriptscriptstyle\text{\rm dyn}}\,\xi_{jt^{\prime}}^{\scriptscriptstyle\text{\rm dyn}}]=0$ for all $i,\,j\in\mathbb{N}$ and $t,\,t^{\prime}\in\mathbb{Z}$ (see Section 3). In our canonical characterization (3.3) of the GDFM decomposition, this orthogonality at all leads and lags is not even an assumption and follows from the definition of $\mathcal{H}^{\bf X}_{\scriptscriptstyle\text{\rm dyn com}}$ and $\mathcal{H}^{\bf X}_{\scriptscriptstyle\text{\rm dyn idio}}$.
On the other hand, a variety of orthogonality assumptions can be found in the literature on the static approximate factor model. [^2] These assumptions, generally, are stated as orthogonality conditions between the factors $f_{t}$ and the idiosyncratic ${\xi}_{it}^{\scriptscriptstyle\text{\rm{stat}}}$ rather than between the common component ${\chi}_{it}^{\scriptscriptstyle\text{\rm{stat}}}$ and the idiosyncratic ${\xi}_{it}^{\scriptscriptstyle\text{\rm{stat}}}$ but, since the loadings are deterministic, the two formulations are equivalent.
  1. (A)
    ${\rm E}[\chi_{it}^{\scriptscriptstyle\text{\rm stat}}\,\xi_{jt}^{\scriptscriptstyle\text{\rm stat}}]=0$, $i,\,j\in\mathbb{N}$ and $t\in\mathbb{Z}.$, i.e., orthogonality between the common and idiosyncratic components only holds contemporaneously. This is assumed in Chamberlain (1983), Chamberlain and Rothschild (1983), De Mol et al. (2008, 2024), Fan et al. (2013), Gersing et al. (2023). And this is the way we characterize the static approximate factor model decomposition (5.2) in Section 5. Then, as we show in Section 5, and except for the assumption that $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ is the linear combination of a finite number $r$ of factors, the static approximate factor model decomposition of $X_{it}$ into $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}+\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$ is not a statistical model, but a representation result that always holds.
  2. (B)
    ${\rm E}[\chi_{it}^{\scriptscriptstyle\text{\rm stat}}\,\xi_{jt^{\prime}}^{\scriptscriptstyle\text{\rm stat}}]=0$, $i,\,j\in\mathbb{N}$ and $t,\,t^{\prime}\in\mathbb{Z}.$, i.e., orthogonality between common and idiosyncratic components at all leads and lags. This is assumed, for instance, in Forni et al. (2009) and Forni and Lippi (2024). Then the static approximate factor model is a genuine statistical model, imposing very strict constraints on the data-generating process. This also makes the static a particular case of the dynamic one (up, perhaps, to some special cases where non-pervasive leads or lags of the factors are orthogonal to their present values).
  3. (C)
    $\{\chi_{it}^{\scriptscriptstyle\text{\rm stat}}|\,i\in\mathbb{N},\,t\in{\mathbb{Z}}\}$ and $\{\xi_{it}^{\scriptscriptstyle\text{\rm stat}}|\,i\in\mathbb{N},\,t\in{\mathbb{Z}}\}$ are mutually independent processes. This, which is assumed in Bai and Ng (2006), Doz et al. (2011), Barigozzi and Luciani (2024), Anderson et al. (2022), and Forni et al. (2023), among others, is a strong reinforcement of (B).
  4. (D)
    $\sup_{n,T\in{\mathbb{N}}}\max_{i,j=1,\ldots,n}{\rm E}[T^{-1}(\sum_{t=1}^{T}{\chi}_{it}^{\scriptscriptstyle\text{\rm{stat}}}{\xi}_{jt}^{\scriptscriptstyle\text{\rm{stat}}})^{2}]\leq M$ for some finite $M>0$. This is assumed by Bai (2003), Doz et al. (2012), and Bai and Li (2016), [^3] Note that Stock and Watson (2002a) do not formally require (D) but somehow use it in their proofs. who call it an assumption of weak dependence between common and idiosyncratic components. It requires finite fourth-order moments, something (B) does not, and is used for proving the consistency of the sample covariance matrix as an estimator of its population counterpart, which in turn is needed to prove the consistency of PCA estimators of the $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$’s. Intuitively close to (B), it is not directly comparable to (A), (B), or (C).
  5. (E)
    Finally, some authors assume deterministic factors (Lawley and Maxwell, 1971; Bai and Li, 2012; Onatski, 2012). This approach is tantamount to conditioning on the factors or, equivalently, on the idiosyncratic components. Apparently, it solves the orthogonality issue since the $\chi_{it}$’s are no longer random. But this is only an apparent solution: conditioning on $\bm{\chi}$, if orthogonality between $\bm{\chi}$ and $\bm{\xi}$ does not hold at all leads and lags, for instance, alters the mean of $\bm{\xi}$—the conditional expectation of $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$, typically, no longer is zero. Treating factors as deterministic quantities while maintaining the assumption that the $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$’s have expectation zero, thus, kind of brings back through the window an orthogonality assumption that was thrown out the door.
As we shall see in Section 9, these orthogonality assumptions (or their absence) play a major role in the problem of combining forecasts of the $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$’s and the $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$’s into forecasts of the $X_{it}$’s.
5 The Static Approximate Factor Model revisited
@@
Paralleling the definition (3.3) of the GDFM decomposition in the previous section, a representation-based characterization of the static approximate factor decomposition of Chamberlain and Rothschild (1983) also can be obtained, which extends (2.1).
@@
Throughout, let us assume that the process $\bf X$ is second-order stationary. Denoting by ${\mathcal{H}}^{{\bf X}_{t}}$ the Hilbert space spanned (at time $t$) by ${\bf X}_{t}$, consider the following decomposition of ${\mathcal{H}}^{{\bf X}_{t}}$ into a statically (at time $t$) common subspace ${\mathcal{H}}^{{\bf X}_{t}}_{\scriptscriptstyle\text{\rm stat com}}$ and its orthogonal complement (still at time $t$, and with respect to ${\mathcal{H}}^{{\bf X}_{t}}$) ${\mathcal{H}}^{{\bf X}_{t}}_{\scriptscriptstyle\text{\rm stat idio}}\coloneqq({\mathcal{H}}^{{\bf X}_{t}}_{\scriptscriptstyle\text{\rm stat com}})^{\perp}$.
@@
A random variable $\zeta$ with values in ${\mathcal{H}}^{{\bf X}_{t}}$ and variance $\sigma^{2}_{\zeta}$ is called statically common-at-time-$t$ if either $\sigma^{2}_{\zeta}=0$ (hence $\zeta=0$ a.s.) or $\sigma^{2}_{\zeta}>0$ and $\zeta/\sigma_{\zeta}$ is the limit in quadratic mean, as $n\to\infty$, of a sequence of standardized elements of ${\mathcal{H}}^{{\bf X}_{t}}$ of the form ${w_{\bf X}^{(n)}}/\sigma^{(n)}_{w}$, where
@@
The spaces $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ and $\mathcal{H}_{\scriptscriptstyle\text{\rm stat idio}}^{{\bf X}_{t}}$ always exist and are unique for all $t$. Intuitively, $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ is the Hilbert space generated by the exploding (as $n\to\infty$) normed linear combinations of the $X_{it}$’s, $i=1,\ldots,n$, for given $t$. Contrary to its dynamic counterpart $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{\bf X}$, it does depend on $t$. It readily follows from the definitions that, for all $t\in\mathbb{Z}$, $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}\subseteq\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{\bf X}$.
@@
where $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$ are the projections of $X_{it}$ on $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ and $\mathcal{H}_{\scriptscriptstyle\text{\rm stat idio}}^{{\bf X}_{t}}$, respectively. In contrast with the dynamic decomposition (3.3), however, orthogonality between the common-at-time-$t$ and the idiosyncratic-at-time-$t$ components $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$ only holds at time $t$—more precisely, $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{i^{\prime}t}^{\scriptscriptstyle\text{\rm stat}}$ are mutually orthogonal for all $i,\,i^{\prime}$, and $t$—while for $\chi_{it}^{\scriptscriptstyle\text{\rm dyn}}$ and $\xi_{it}^{\scriptscriptstyle\text{\rm dyn}}$ in (3.3) and (3.5) orthogonality of $\chi_{it}^{\scriptscriptstyle\text{\rm dyn}}$ and $\xi_{i^{\prime}t^{\prime}}^{\scriptscriptstyle\text{\rm dyn}}$ holds for all $i,\,i^{\prime}$, $t$, and $t^{\prime}$.
Alone, (5.2) does not imply (2.1) along with conditions (a)–(f), though. Adding the requirement that the eigenvalues
@@
of the $n\times n$ covariance matrices ${\bm{\Gamma}}^{(n)}_{{X}}$ of ${\bf X}^{(n)}_{t}\coloneqq(X_{1t},\ldots,X_{nt})^{\prime}$ (which, in view of second-order stationarity, do not depend on $t$) are such that
$\lim_{n\to\infty}\lambda^{(n)}_{X;r}=\infty\quad\text{ and }\quad\lim_{n\to\infty}\lambda^{(n)}_{X;r+1}<\infty$ (5.3)
@@
for some finite $r\in{\mathbb{N}}$ independent of $n$, it can be shown (see Proposition 1 below) that $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ admits $r$-dimensional orthonormal bases ${\bf f}_{t}=(f_{1t},\ldots,f_{rt})^{\prime}$, any of which can be interpreted as the value at time $t$ of an $r$-tuple of factors, yielding a Chamberlain and Rothschild static approximate factor decomposition (2.1) satisfying the assumptions (a)–(f) of Section 2.1, with statically common component $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ of the form ${\bf B}_{i}{\bf f}_{t}$.
@@
The process $\bf X$ being second-order stationary, it admits the canonical representation (5.2). Despite the fact that the Hilbert spaces  $\mathcal{H}^{{\bf X}_{t}}$, $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$, and $\mathcal{H}_{\scriptscriptstyle\text{\rm stat idio}}^{{\bf X}_{t}}$ depend on $t$, its $n\times n$ covariance matrices ${\bm{\Gamma}}^{(n)}_{X}$, their eigenvalues and the corresponding eigenvectors, do not; accordingly, the number $r$ of diverging eigenvalues in (5.3), their values and the corresponding eigenvectors, and the projection operators from $\mathcal{H}^{{\bf X}_{t}}$ to $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ and $\mathcal{H}_{\scriptscriptstyle\text{\rm stat idio}}^{{\bf X}_{t}}$ (which only depend on ${\bm{\Gamma}}^{(n)}_{X}$), are the same for all $t$.
The following result shows that (5.3) yields a characterization of the static approximate factor model which is much more parsimonious than the Chamberlain and Rothschild one; see the Appendix for a proof.
Proposition 1.
Let $\bf X$ be second-order stationary. Then, the following statements are equivalent:
  1. (A)
    (5.3) holds, that is, $\lim_{n\to\infty}\lambda^{(n)}_{X;r}=\infty$ and $\lim_{n\to\infty}\lambda^{(n)}_{X;r+1}<\infty$;
  2. (B)
    (2.1) holds and satisfies (a)–(f), that is, $\bf X$ admits a static approximate factor model representation with $r$ factors in the sense of Section 2.
Moreover, if (A)—hence also (B)—holds, then (2.1) coincides with (5.2): namely, $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$ in (2.1) and $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$ in (5.2) coincide for all $i\in\mathbb{N}$ and $t\in\mathbb{Z}$.
In other words, the static approximate factor model with $r$ factors as defined in Section 2 (a definition that, unlike the Chamberlain and Rothschild one, includes (d2)) is entirely characterized by (A), i.e., by the divergence of $\lambda^{(n)}_{X;r}$ and the boundedness of $\lambda^{(n)}_{X;r+1}$ as $n\to\infty$—that is, by the existence of a growing gap, in the spectrum of the covariance matrices ${\bm{\Gamma}}^{(n)}_{X}$ of $\bf X$, between the $r$th and $(r+1)$th eigenvalues.
The advantage of characterization (A) over characterization (B) is not only its parsimony but, above all, the fact that it only involves features of the observable process $\bf X$ while (B) is about the unobservable common and idiosyncratic components of an unobservable decomposition (2.1). The desire for an observable-based characterization of their model is also very much present in Chamberlain and Rothschild (1983), all the more so that the unobservable decomposition (2.1), in their approach, is a postulated one. In their Theorem 4, they make a step into the direction of such an observable-based characterization by showing that (5.3) implies the decomposition of ${\bm{\Gamma}}^{(n)}_{X}$ (for any $n\geq r)$ into the sum of a covariance matrix with rank $r$ and a matrix with uniformly bounded (as $n\to\infty$) eigenvalues, hence that (A) implies their definition (which requires (d1) but not (d2): see their Definition 2) of a static approximate factor model. However, because (d2) is not included in that definition, the converse does not hold, and that definition does not imply the divergence of $\lambda^{(n)}_{X;r}$, hence cannot imply (A). Therefore, Chamberlain and Rothschild (1983) cannot state an equivalence result similar to Theorem 1.
This is, of course, a very minor logical issue, and adding (d2) to the Chamberlain and Rothschild definition immediately restores its equivalence with (A). Nevertheless, and since their proof of Theorem 4 is scattered over several sections and involves some technical terminology specific of financial econometrics, we are providing, in the Appendix, an independent and purely mathematical proof of Theorem 1.
@@
Finally, it is important to note that neither (5.3) nor Chamberlain and Rothschild (1983) impose any rate on the divergence, as $n\to\infty$, of the eigenvalues of the covariance matrices ${\bm{\Gamma}}^{(n)}_{\chi}$ and ${\bm{\Gamma}}^{(n)}_{X}$, which in principle can be sublinear, linear, or superlinear. Indeed, no divergence rate assumption is needed for identifying the model as long as an eigen-gap widening, as $n\to\infty$, to infinity exists between bounded and unbounded eigenvalues. When turning to estimation, however, all works mentioned above (Stock and Watson, 2002a, Bai, 2003, Forni et al., 2009, etc.) make the additional assumption of a linear rate of divergence of the unbounded eigenvalues, and the same linear divergence assumption is made by Bai and Ng (2002) when proposing an information criterion to determine $r$ based on the behavior of sample eigenvalues (see Section 8). That linear divergence assumption is required for the consistency of the estimation methods, though, not by the identifiability of the number of factors. We argue, in Section 7, that such linear divergence automatically follows under the quite natural additional assumption of cross-sectional exchangeability.